OpenStudy (anonymous):

Need help please what is 67% of 700?

OpenStudy (anonymous):

what are the steps to getting this ?

OpenStudy (theviper):

x% of y is\[\LARGE{\frac{x}{100}\times y \Huge{\color{red}{\checkmark}}}\]

OpenStudy (theviper):

So 67% of 700 is\[\LARGE{\frac{67}{100}\times700=??}\]

OpenStudy (theviper):

Got it??

OpenStudy (anonymous):

so I just need to divide 67 by 100 then times that ? by 700

OpenStudy (anonymous):

so its 469

OpenStudy (anonymous):

is that right ?

OpenStudy (theviper):

Yes u got it dude :D

OpenStudy (anonymous):

so I would say in my answer 469 percent ?

OpenStudy (theviper):

no its a number only 469 is 67% of 700 :)

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

will do

OpenStudy (anonymous):

it says your offline

OpenStudy (theviper):

nope but u can be my fan :)

OpenStudy (anonymous):

I clicked on become a fan its not letting me ?

OpenStudy (theviper):

me too :(